OpenStudy (anonymous):

Mary reduced the size of a painting to a width of 3.3 in. What is the new height if it was originally 32.5 in tall and 42.9 in wide?

OpenStudy (anonymous):

32.5/42.9 = x/3.3 42.9x =32.5*3.3 x = (32.5*3.3)/42.9 =2.5

OpenStudy (anonymous):

word

OpenStudy (anonymous):

peace!